Elements of

Abstract and Linear Algebra

E. H. Connell

E.H. Connell Department of Mathematics University of Miami P.O. Box 249085 Coral Gables, Florida 33124 USA ec@math.miami.edu

Mathematical Subject Classifications (1991): 12-01, 13-01, 15-01, 16-01, 20-01

\begin{matrix}c \\ ◯ \end{matrix}1999 E.H. Connell

March 20, 2004